OpenStudy (anonymous):

Given f(x) = the quantity of 4x plus 1, divided by 3, solve for f-1(3).

OpenStudy (anonymous):

f(x) = (4x +1)/3 so f^-1(x) = (3x - 1)/4 f^-1(3) = (3(3) -1)/4 = (9-1)/4 = 8/4 = 2